# GIGA: Generalizable Sparse Image-driven Gaussian Avatars

### v1

@@ 3.3 Training
GIGA generates Gaussian texture maps at $512\times 512$ resolution. As some texels in these textures are not associated with any triangle, they are discarded during rendering.
@@
Reconstruction Term. Following Kerbl et al. [13], we compute the mean absolute error $\mathcal{L}_{\mathrm{L1}}$ and the structural similarity measure $\mathcal{L}_{\mathrm{ssim}}$ [53] between the rendered image $I_{k}$ and the ground-truth image $I_{\mathrm{gt},k}$. We additionally evaluate the VGG-based [43] perceptual loss [61] $\mathcal{L}_{\mathrm{LPIPS}}$ on randomly sampled patches with centers within the ground truth segmentation mask $A_{\mathrm{gt},k}$. To ensure scale-invariance of the perceptual features, we sample 16 patches of varying sizes: $128\times 128$, $256\times 256$, and $512\times 512$; then resize all patches to $256\times 256$ before computing the loss, following Cao et al. [2]. To improve outlines of Gaussian primitives, we also compute $\mathcal{L}_{\mathrm{mask}}$, the mean squared error between the rendered opacity images $A_{k}$ and ground truth character segmentation masks $A_{\mathrm{gt},k}$. The final reconstruction term is averaged over all images in the batch and defined as:
@@
Due to UV mapping distortions, some mesh regions may be oversampled with Gaussians. The poor initial model state will encourage some Gaussians to become transparent and inadequately large and move away from the mesh. To reclaim those Gaussians for actual modeling, we additionally penalize offsets of low-opacity primitives with $\mathcal{L}_{\delta\mathbf{x}_{0};\alpha}$, where the indicator function $\mathds{1}_{\left[\alpha<\epsilon\right]}$ triggers the extra penalty only when Gaussian opacity $\alpha$ is below a threshold $\epsilon$. Importantly, we disable gradient flow through the indicator function. We similarly penalize scales $\mathbf{s}^{\prime}$ when opacity is below $\varepsilon$ with $\mathcal{L}_{\mathbf{s}^{\prime};\alpha}$. Following Lombardi et al. [22], we encourage all Gaussians to be either completely opaque or completely transparent by computing negative log-likelihood of the beta distribution $\textrm{Beta}\left(0.5,0.5\right)$ as $\mathcal{L}_{\alpha}$.
@@ 3.4 Implementation Details
We train GIGA for $500$k iterations on a single NVIDIA H100 GPU using 1 subject per batch. All models are optimized with AdamW [25], with the learning rate being linearly increased from 00 to $1e^{-4}$ during the first $25$k training steps. We also set the weight decay parameter to $1e^{-4}$.
@@
We render $4$ random views for each input subject in the dataset image resolution to compute the training objective.
Dataset MVHumanNet THuman2.0
Method     Metric PSNR SSIM LPIPS PSNR LPIPS
NHP 17.72 0.5678 402.7
@@
GHG 21.90 133.4
GIGA 22.19 0.7526 70.2 24.70 51.6
Table 1: Quantitative Comparison. We evaluate generalization to unseen identities of Neural Human Performer (NHP) and TransHuman (TH) on MVHumanNet, and Generalizable Human Gaussians (GHG) on THuman2.0.
@@ 4 Results
Table 2: Ablation Results for GIGA. We examine our design choices, proposed in Sec. 3, see detailed analysis in Sec. 4.4.
We first explain the datasets and metrics (Sec. 4.1). Then, we show qualitative results (Sec. 4.2) and comparisons (Sec. 4.3). Lastly, we ablate our design choices (Sec. 4.4).
4.1 Dataset and Metrics
Datasets. To evaluate GIGA, we use the multi-view human performance capture datasets MVHumanNet [55] and DNA-Rendering [3], and also static 3D human scans dataset from THuman2.0 [59]. We split MVHumanNet as follows: IDs $100000-100999$ are used for training and IDs $101000-102386$ are held out for validation and testing. From the testing part, we choose $40$ subjects for quantitative evaluation. For experiments with DNA-Rendering, we use the released Part-2 for training (400 subjects) and Part-1 for validation (40 subjects). In both cases, we select $12$ cameras for supervision by picking every $4$th camera from the provided camera calibration. From each subject sequence, we sample every $20$th pose for training, but not more than $15$ poses per character. We follow the GHG [16] training and validation split of THuman2.0 and render multi-view images from original 3D scans similarly to GHG.
4.2 Qualitative Results
Fig. 34, and 6 (and the supplementary video) present GIGA’s novel view rendering results of unseen subjects performing unseen poses. Notably, GIGA achieves photorealistic and view-consistent rendering and effectively captures fine details such as clothing wrinkles and intricate textures. We highlight that for novel subjects, GIGA achieves a rendering quality comparable to that of the training subjects, demonstrating its generalization ability to novel identities.
4.3 Comparison
Competing Methods. We compare GIGA to other generalizable dynamic image-driven methods discussed in Sec. 2: NHP [15] and TransHuman (TH) [30]. We exclude NNA [51] from comparisons, as we could not reproduce the training results from the code provided by the authors. We train NHP and TH using the same training/validation split of MVHumanNet and follow the same training setup as for our method. We also adopt GHG [16] as a baseline and train GIGA on the THuman2.0 dataset [59].
@@ 4.3 Comparison
Training set $\to$ Testing set
Metric DNA $\to$ DNA MVH $\to$ DNA
$\uparrow$ PSNR 19.63 19.49
$\uparrow$ SSIM 0.7255 0.7138
$\downarrow$ LPIPS 106.4 114.9
Baselines. Tab. 1 provides quantitative comparisons of GIGA against state-of-the-art baselines for generalizable human rendering. In all cases, we evaluate the generalization ability to unseen identities from held-out validation sequences. Fig. 3 provides an overview of qualitative results. NHP relies on sparse 3D convolutions to process volumetric features in the observed pose space, thus, suffering from missing input signals due to occlusions and failing to generalize to unseen identities. Despite operating in a canonical template pose space and tokenizing the template for processing with a transformer-based network, TransHuman also cannot learn meaningful priors from the large data collection. Both NHP and TransHuman are extremely slow due to implicit representations at their cores, which significantly limits their generalization abilities. GIGA, on the other hand, utilizes the power of the shared texel space to a maximum degree: all feature representations for digital humans are defined in the same texel space, and intermediate features enhance the quality of the final prediction through skip-connections. MultiHeadUNet is also significantly more computationally efficient, which explains, both, qualitative and quantitative improvements over baselines.
While being much more efficient than previous works, GHG targets only static reconstruction from sparse input views. GHG also models humans as a set of 3D Gaussian scaffolds in the observed pose space and cannot be easily extended to dynamic scenarios. GHG handles Gaussian color prediction by pretraining a separate texture inpainting network. GIGA learns to operate with appearance and geometry features simultaneously, yielding results of higher quality, both, visually (Fig. 4) and quantitatively (Tab. 1).
Cross-Dataset Validation. To demonstrate cross dataset generalization we also train two variants of GIGA where one is trained and tested on the DNA-Rendering dataset while the other variant is trained on MVHumanNet and tested on the DNA-Rendering. We follow train/test splits specified in Sec. 4.1. Tab. 4 and Fig. 6 demonstrate that our method is capable to generalize across datasets as the model trained on the large-scale MVHumanNet data performs comparable on DNA-Rendering, which clearly shows that our method effectively learns the prior from large datasets.
4.4 Ablation Study
Model Architecture. To benchmark MultiHeadUNet (IV) at the core of GIGA, we propose 3 alternative architectures with approximately the same number of trainable parameters ($\simeq$90M): a simple encoder-decoder model (I), a conventional UNet with skip-connections between corresponding up- and downsampling blocks of the encoder and decoder (II), and a model with 2 encoders $\mathcal{E}_{\mathrm{a}},\mathcal{E}_{\mathrm{g}}$ and 3 decoders $\mathcal{D}_{\mathrm{a}},\mathcal{D}_{\mathrm{p}},\mathcal{D}_{\mathrm{g}}$, but without skip-connections (III). We observe, both, quantitatively (Tab. 2) and qualitatively (Fig. 5) that the configuration IV leads to a higher quality overall, being particularly helpful at preservation of fine appearance details, observed in the input signal.
[…]
Motion Conditioning. Configurations I-IV use an MLP-based motion encoder $\mathcal{E}_{\mathrm{m}}$ and cross-attention motion conditioning by default. We additionally remove cross attention blocks from IV, leaving intermediate features $\mathbf{F}_{\mathrm{uv}}^{\mathrm{a}},\mathbf{F}_{\mathrm{uv}}^{\mathrm{g}}$ without motion conditioning (IVa). As it is evident from the quantitative evaluation (2), enabling motion conditioning via cross-attention improves the visual quality of GIGA’s Gaussian avatars.
@@ 4.4 Ablation Study
Scaling Dataset Size. GIGA also benefits from training on large data collections, as reported in Tab. 3 (please see the supplemental materials for the qualitative comparisons). While scaling a digital human reconstruction model to hundreds of identities presents a challenge as evidenced by the fact that prior works failed to train on such large training corpora, it leads to an improved generalization overall.
@@
Original skeleton tracking and SMPL-X parameters in MVHumanNet are obtained from a dense set of views. To mimic a real-world scenario, we re-track a small number of sequences from a set of 4 views and fit SMPL-X models to estimated 3D keypoints. Qualitative results of GIGA with sparsely tracked SMPL-X template are shown in Fig. 7. If sparse tracking is mostly correctly aligned with the actual human actor, then GIGA can produce results, similar to dense view tracking case. For real-world use cases, a suitable sparse tracking solution should be applied, which is beyond the scope of this paper.
5 Limitations
GIGA shows unprecedented scalability that enables training on thousands of multi-view videos, thanks to our efficient representation and highly scalable architecture, and respective generalization without sacrificing rendering quality. However, it still faces some limitations that should be addressed in the future. While using SMPL-X as body template greatly facilitates generalization, it does not allow us to handle non-rigid dynamics (e.g., hair and loose clothing) properly without additional assumptions or physics-based priors. Here, a more advanced human shape prior that includes clothing geometry might alleviate some of these limitations [60, 4, 38]. Moreover, the dependency on parametric body models and motion tracking leads to quality degradation in case of template misalignment or inaccurate tracking. Future work could explore end-to-end optimization of body shape and pose parameters, which has already proven to be successful for face-only rendering approaches [46].
6 Conclusion
This work presented GIGA, a generalizable sparse image-drivable Gaussian avatar. Trained on a large-scale multi-view dataset, GIGA synthesizes texel-aligned 3D Gaussian avatars from sparse input views in a feed-forward manner. Our approach achieves state-of-the-art generalization to unseen identities while preserving person-specific pose-dependent appearance changes thanks to our scalable architecture and efficient representation. We believe our proposed model could benefit future research in this domain and take another step towards enabling more accessible and immersive remote collaboration.
Supplementary Material
7 Projected Texture Estimation
Partial Texture Computation. After articulating the normalized canonical mesh to the observed pose $\boldsymbol{\theta}_{j}$, we obtain posed vertex coordinates $\mathbf{x}_{j}$. The next step is an initialization of texel coordinates buffer $\mathbf{T}_{\mathbf{x}}\in\mathbb{R}^{T\times T\times 3}$. We set $\mathbf{x}_{j}$ as attributes to the mesh $\mathcal{V}(\boldsymbol{\theta}_{j},\boldsymbol{\beta})$ vertices and perform texture sampling w.r.t UV parametrization $\mathcal{M}_{\mathrm{uv}}$ to fill buffer $\mathbf{T}_{\mathbf{x}}$ with coordinates of posed texels. In the following, we will drop the pose index $j$, assuming that all operations are performed for the observed pose. For each input view $k$, pixel coordinates are calculated for each texel:
Visibility Check and Texture Aggregation.
Figure 8: Qualitative Results from Front View. More qualitative results produced by GIGA. All subjects are from the test split. GT stands for ground truth.
Not every texel is observed from the view $\pi_{k}$. Hence, we need to discard invisible texels from the partial texture. We first render a depth image of the body template $\mathcal{V}\left(\boldsymbol{\theta},\boldsymbol{\beta}\right)$ and retrieve vertex visibility buffer provided by the differentiable rasterizer [17]. After barycentric interpolating the visibility buffer $\mathcal{M}_{\mathrm{uv}}$, we obtain the visibility mask $\mathbf{V}_{\mathrm{uv},k}\in\mathbb{R}^{T\times T}$. Then, we compute angle visibility scores $\mathbf{V}^{a}_{\mathrm{uv},k}\in\mathbb{R}^{T\times T}$:
[…]
8 Baselines
We use the open-source implementations of both NHP [15] and TransHuman [30] for training. For the sake of efficiency, we replace original attention layers in baselines with their more efficient analogs [18]. Moreover, TransHuman requires clusterization of body template vertices. Therefore, we follow the original codebase and cluster SMPL-X vertices using K-Means with $K=300$ clusters. For evaluation against GHG [16], we use validation data released by authors. GHG metric numbers are taken from the original paper.
9 More Qualitative Results
Here we show additional qualitative examples from GIGA results. Fig. 8 presents additional qualitative results for subjects from the testing split. We also present free-viewpoint renderings in Fig. 10 for subjects used neither for training nor testing.
Figure 10: Free-viewpoint Rendering. Views from a circular trajectory around unseen characters.

### v2

@@ 3.3 Training
GIGA generates texel-aligned 3D Gaussian texture maps at $512\times 512$ resolution. As some texels in these textures are not associated with any triangle, they are assigned zero opacity during rendering.
@@
Reconstruction Term. Following Kerbl et al. [15], we compute the mean absolute error $\mathcal{L}_{\mathrm{L1}}$ and the structural similarity measure $\mathcal{L}_{\mathrm{ssim}}$ [59] between the rendered image $I_{k}$ and the ground-truth image $I_{\mathrm{gt},k}$. We additionally evaluate the VGG-based [49] perceptual loss [68] $\mathcal{L}_{\mathrm{LPIPS}}$ on randomly sampled patches with centers within the ground truth segmentation mask $A_{\mathrm{gt},k}$. To improve outlines of Gaussian primitives, we also compute $\mathcal{L}_{\mathrm{mask}}$, the mean squared error between the rendered opacity images $A_{k}$ and ground truth character segmentation masks $A_{\mathrm{gt},k}$. The final reconstruction term is averaged over all images in the batch and defined as:
@@
Figure 4: Qualitative Comparison against 3D Gaussian-based methods. GIGA achieves higher rendering quality than previous GHG [19] and GPS-Gaussian [69] with fewer artifacts. Results reported on THuman2.1 dataset [66].
@@
Due to UV mapping distortions, some mesh regions may be oversampled with Gaussians. The poor initial model state will encourage some Gaussians to become transparent and inadequately large and move away from the mesh. To reclaim those Gaussians for actual modeling, we additionally penalize offsets of low-opacity primitives with $\mathcal{L}_{\delta\mathbf{x}_{0};\alpha}$, where the indicator function $\mathds{1}_{\left[\alpha<\epsilon\right]}$ triggers the extra penalty only when Gaussian opacity $\alpha$ is below a threshold $\epsilon$. Importantly, we disable gradient flow through the indicator function. We similarly penalize scales $\mathbf{s}$ when opacity is below $\varepsilon$ with $\mathcal{L}_{\mathbf{s}^{\prime};\alpha}$. Following Lombardi et al. [26], we encourage all Gaussians to be either completely opaque or completely transparent by computing the negative log-likelihood of the beta distribution $\textrm{Beta}\left(0.5,0.5\right)$ as $\mathcal{L}_{\alpha}$.
@@ 3.4 Implementation Details
We train GIGA for 250250k iterations on a single NVIDIA H100 GPU using 4 subjects per batch. All models are optimized with AdamW [29], with the learning rate being linearly increased from 0 to $1e^{-4}$ during the first 2525k training steps. We also set the weight decay parameter to $1e^{-4}$.
@@
Dataset Method $\uparrow$ PSNR $\uparrow$ SSIM $\downarrow$ LPIPS
MVHumanNet NHP 17.72 0.5678 402.7
@@
GIGA 22.19 0.7526 70.2
THuman2.1 GPS-G 18.22 0.7768 207.9
GHG 18.86 0.7394 76.5
GIGA 20.19 0.7944 58.7
Table 1: Quantitative Comparison. We evaluate generalization to unseen identities of Neural Human Performer (NHP) and TransHuman (TH) on MVHumanNet; GPS-Gaussian (GPS-G) and Generalizable Human Gaussians (GHG) on THuman2.1.
@@ 4 Results
We first explain the datasets and metrics (Sec. 4.1). Then, we show qualitative results (Sec. 4.2) and comparisons (Sec. 4.3). Lastly, we ablate our design choices (Sec. 4.4). Supplementary materials contain additional qualitative examples.
4.1 Dataset and Metrics
Datasets. To evaluate GIGA, we use the multi-view human performance capture datasets MVHumanNet [61] and DNA-Rendering [3], and also the 3D human scan dataset THuman2.1 [66]. Details on the data usage protocol for training can be found in the supplementary material.
4.2 Qualitative Results
Fig. 34, and 5 (and the supplementary video) present GIGA’s novel view rendering results of unseen subjects performing unseen poses. Notably, GIGA achieves photorealistic and view-consistent rendering and effectively captures fine details such as clothing wrinkles and intricate textures. We highlight that for novel subjects, GIGA achieves a rendering quality comparable to that of the training subjects, demonstrating its generalization ability to novel identities. Please see the in-the-wild examples from Neuman dataset in the supplementary materials.
Training set $\to$ Testing set
Metric DNA $\to$ DNA MVH $\to$ DNA
$\uparrow$ PSNR 19.63 19.49
$\uparrow$ SSIM 0.7255 0.7138
$\downarrow$ LPIPS 106.4 114.9
4.3 Comparison
Competing Methods. We compare GIGA to other generalizable dynamic image-driven methods discussed in Sec. 2, i.e., NHP [18], TransHuman (TH) [34], GPS-Gaussian [69], and GHG [19]. We exclude concurrent human avatar reconstruction methods IDOL [72], SIGMAN [64], and MoGa [74] from comparison, since they focus on reconstructing animatable avatars from single images.
Tab. 1 provides quantitative comparisons of GIGA against state-of-the-art baselines for generalizable human rendering. In all cases, we evaluate the generalization ability to unseen identities from held-out validation sequences. Fig. 3 provides an overview of qualitative results. NHP relies on sparse 3D convolutions to process volumetric features in the observed pose space and suffers from missing input signals due to occlusions, ultimately failing to generalize to unseen identities. Despite operating in a canonical template pose space and tokenizing the template for processing with a transformer-based network, TransHuman also cannot learn meaningful priors from the large data collection. GIGA, on the other hand, utilizes the power of the shared texel space to a maximum degree: all feature representations for digital humans are defined in the same texel space, and intermediate features enhance the quality of the final prediction through skip-connections. MultiHeadUNet is also significantly more computationally efficient, which explains both qualitative and quantitative improvements over baselines.
GPS-Gaussian heavily relies on depth estimation from two adjacent source views to unproject pixel-aligned 3D Gaussians for novel view rendering. As such, it is limited by the quality of the depth estimation, which might fail if source views are too far from each other. Artifacts of erroneous depth-based unprojection are visible in Fig. 4.
GHG targets only static reconstruction from sparse input views; it models humans as a set of 3D Gaussian scaffolds in the observed pose space and cannot be easily extended to dynamic scenarios. GHG handles Gaussian color prediction by pretraining a separate texture inpainting network. GIGA learns to operate with appearance and geometry features simultaneously, yielding results of higher quality, both qualitatively (Fig. 4) and quantitatively (Tab. 1).
Cross-Dataset Validation.
[…]
To demonstrate the cross-dataset generalization ability, we train two variants of GIGA, where one is trained and tested on the DNA-Rendering [61] dataset, while the other variant is trained on MVHumanNet [3] and tested on the DNA-Rendering dataset. As seen in Tab. 2 and Fig. 5, our method is capable of generalizing across datasets with minimal quality drop, as the model trained on the MVHumanNet dataset performs comparably to the counterpart trained on the DNA-Rendering dataset, which clearly shows that our method effectively learns the prior from large datasets.
Single/Sparse-View Inputs. To enable a single-input-view regime at test-time, for each character we precompute a static RGB texture from multiple views and concatenate it with $\mathbf{T}_{\mathbf{x}_{0}}$ from Eq. (2) along channel dimension. This serves as ”mean” appearance information to be used for unobserved parts of the template by the network. We retrain GIGA with this additional input and evaluate the effect of the number of views at test-time in Tab. 3. The 4-view setup (front, back, left, and right; 0c) provides a full coverage of the character, whereas the 2-view setup (0b) only uses the front and the back views, resulting in slightly better metric values. In the single view setup (0a), we only provide the frontal view, which contains information about the dynamic appearance, while the backside of the character will be reconstructed mainly based on the precomputed static texture - thus leading to reduced quantitative performance. We report qualitative results in the supplementary material.
4.4 Ablation Study
In this section, we discuss our design choices for the architecture and training objective of GIGA. We also examine the effects of different sizes of training datasets and the influence of motion tracking from sparse input views in the supplementary material.
Model Architecture. To benchmark MultiHeadUNet (IV) at the core of GIGA, we propose 3 alternative architectures with approximately the same number of trainable parameters ($\simeq$90M): a simple encoder-decoder model (I), a conventional UNet with skip-connections between corresponding up- and downsampling blocks of the encoder and decoder (II), and a model with 2 encoders $\mathcal{E}_{\mathrm{a}},\mathcal{E}_{\mathrm{g}}$ and 3 decoders $\mathcal{D}_{\mathrm{a}},\mathcal{D}_{\mathrm{p}},\mathcal{D}_{\mathrm{g}}$, but without skip-connections (III). We observe, both quantitatively (Tab. 3) and qualitatively (Fig. 6), that the configuration IV leads to a higher quality overall, being particularly helpful at preservation of fine appearance details, observed in the input signal.
Motion Conditioning. Configurations I-IV use an MLP-based motion encoder $\mathcal{E}_{\mathrm{m}}$ and cross-attention motion conditioning by default. We additionally remove cross attention blocks from IV, leaving intermediate features $\mathbf{F}_{\mathrm{uv}}^{\mathrm{a}},\mathbf{F}_{\mathrm{uv}}^{\mathrm{g}}$ without motion conditioning (IVa). As it is evident from the quantitative evaluation (3), enabling motion conditioning via cross-attention improves the visual quality of GIGA’s predictions.
@@ 4.4 Ablation Study
Table 3: Ablation Results for GIGA. We examine our design choices, proposed in Sec. 3, see detailed analysis in Sec. $4.$
5 Limitations
GIGA shows scalability to hundreds of multi-view videos, thanks to our efficient representation and efficient architecture, and respective generalization without sacrificing rendering quality. Future work could explore the construction of a latent space for virtual humans to improve temporal consistency. SMPL-X as a body template forms the basis for generalization, but it does not allow for handling non-rigid dynamics (e.g., hair and loose clothing) without additional assumptions or physics-based priors. More advanced human shape prior that includes clothing geometry might alleviate some of these limitations [67, 4, 43]. To mitigate motion tracking issues, an end-to-end optimization of body shape and pose parameters could be a promising next step, which has already proven to be successful for face-only rendering approaches [52].
6 Conclusion
This work presented GIGA, a generalizable, sparse image-driven 3D Gaussian human model. Trained on a large-scale multi-view dataset, GIGA dynamically synthesizes texel-aligned 3D Gaussians from sparse input videos in a feed-forward manner. Our approach achieves state-of-the-art generalization to unseen identities while preserving person-specific pose-dependent appearance changes thanks to our scalable architecture and efficient representation. We believe our proposed model could benefit future research in this domain and take another step towards enabling more accessible and immersive remote collaboration.
Supplementary Material
In this supplementary document, we first describe the process of input texture estimation in Sec. A. Second, in Sec. B we provide details about training of GIGA, specifying data splits (B.1) and the training protocol (B.2). Third, we present training and evaluation setups for the baselines in Sec. C. Then, in Sec. D we elaborate on the test time regime of GIGA with only a monocular input. Afterwards, we report in-the-wild examples on a monocular videos dataset in Sec. E. Finally, we present additional ablation studies conducted for GIGA in Sec. F: the effect of scaling the training dataset size and qualitative results with motion tracking from sparse views.
We also report qualitative results of GIGA in the supplementary video.
Appendix A Projected Texture Estimation.
Partial Texture Computation. After articulating the normalized canonical mesh to the observed pose $\boldsymbol{\theta}_{j}$, we obtain posed vertex coordinates $\mathbf{x}_{j}$. The next step is an initialization of the texel coordinates buffer $\mathbf{T}_{\mathbf{x}}\in\mathbb{R}^{T\times T\times 3}$. We set $\mathbf{x}_{j}$ as attributes to the mesh $\mathcal{V}(\boldsymbol{\theta}_{j},\boldsymbol{\beta})$ vertices and perform texture sampling w.r.t UV parametrization $\mathcal{M}_{\mathrm{uv}}$ to fill buffer $\mathbf{T}_{\mathbf{x}}$ with coordinates of posed texels. In the following, we will drop the pose index $j$, assuming that all operations are performed for the observed pose. For each input view $k$, pixel coordinates are calculated for each texel:
Visibility Check and Texture Aggregation. Not every texel is observed from the view $\pi_{k}$. Hence, we need to discard invisible texels from the partial texture. We first render a depth image of the body template $\mathcal{V}\left(\boldsymbol{\theta},\boldsymbol{\beta}\right)$ and retrieve the vertex visibility buffer provided by the differentiable rasterizer [20]. After barycentric interpolating the visibility buffer $\mathcal{M}_{\mathrm{uv}}$, we obtain the visibility mask $\mathbf{V}_{\mathrm{uv},k}\in\mathbb{R}^{T\times T}$. Then, we compute angle visibility scores $\mathbf{V}^{a}_{\mathrm{uv},k}\in\mathbb{R}^{T\times T}$:
where $\mathbf{N}\in\mathbb{R}^{T\times T\times 3}$ are per-texel normals obtained through barycentric interpolation of the vertex normals. $\operatorname{unit}$ denotes L2-normalization for per-texel viewing direction, and $\left(\odot\right)$ denotes dot product between vectors. Next, we calculate indices $\overline{k}$ of partial texture with highest visibility scores
$\overline{k}=\underset{k}{\operatorname{argsort}}\left(\mathbf{V}^{a}_{\mathrm{uv},k}\right).$ (12)
Appendix B Training Details.
@@ 6 Conclusion
Figure 7: Qualitative Results in Single/Sparse-View Setups. GIGA can synthesize virtual humans from a single or a few (2-4) sparse views.
B.1 Data Splits.
MVHumanNet. We split MVHumanNet as follows: IDs 100000100999100000-100999 are used for training and IDs 101000102386101000-102386 are held out for validation and testing. From each subject sequence, we sample every $20$th pose for training, but not more than $15$ poses per character. From the testing part, we chose $40$ subjects for quantitative evaluation.
DNA-Rendering. For experiments with DNA-Rendering, we use the released Part-2 for training (400 subjects) and Part-1 for validation (39 subjects).
THuman2.1. Since the original release of the THuman2.0 dataset with 526 3D scans, it has been extended with additional 1919 scans and rebranded as THuman2.1. We use the new 1919 scans for training of GIGA and corresponding baselines (GPS-Gaussian [69] and GHG [19]), and use 100 old scans for validation as specified in the GHG code release. We render multi-view images from the scans and use SMPL-X parameters provided by the authors of the dataset.
B.2 Training Procedure.
We render $4$ random views for each input subject to compute the training objective. When computing perceptual loss, we sample 16 patches of varying sizes: $128\times 128$, $256\times 256$, and $512\times 512$; then resize all patches to $256\times 256$ before computing the loss, following Cao et al. [2], to ensure scale-invariance of perceptual features.
Figure 8: In-the-Wild Examples. Qualitative Results on the Neuman dataset.
Appendix C Baselines.
Neural Human Performer and TransHuman. We use the open-source implementations of both NHP [18] and TransHuman [34] for training. For the sake of efficiency, we replace original attention layers in baselines with their more efficient analogs [21]. TransHuman requires clusterization of body template vertices. Therefore, we follow the original codebase and cluster SMPL-X vertices using K-Means with $K=300$ clusters. We train NHP and TransHuman using the same training/validation split of MVHumanNet as described in subsection B.1.
GPS-Gaussian. The publicly released checkpoint of GPS-Gaussian was obtained after training in a 16-view setup. For a fair comparison, we retrain GPS-Gaussian using 6-view setup (60 degrees angular distance between views) - we observed that GPS-Gaussian had failed to converge under sparser view setups. At test-time, our version of GPS-Gaussian works with 2 input views, having at most 67 degrees angular difference, which amounts to approximately 5 views placed around the subject in total.
GHG. GHG contains two models: an inpainting model for incomplete RGB textures and a 3D Gaussian regression model. We use the checkpoint of the inpainting model provided by the authors and then finetune it jointly with the Gaussian regression model. The latter is initialized from the released checkpoint and trained further on a larger collection of 3D scans, as specified in subsection B.1. We follow the training setup of GHG, randomly sampling 3 input views during training. For evaluation, we select 4 views with full coverage of the subject.
Appendix D Single/Sparse Input Views.
As discussed in subsection 5, GIGA can produce results even with a single input video. We report qualitative results in Fig. 7, where we pick the frontal camera view for the ”1 input view” setting, frontal and back cameras for the ”2 input views”; and frontal, back, left, and right cameras with 90 90-degree angular difference for the ”4 input views” setting. However, since GIGA is not trained as a generative model and, hence, lacks generative capabilities, it requires some additional signal to predict meaningful 3D Gaussian parameters in the unobserved body parts.
As it was briefly mentioned in subsection 5, we perform a brief character calibration procedure before inference: if only a single monocular video is available, we select those frames where the various sides of the actor are visible; if multi-view videos are available, we select 4 to 8 views with maximum coverage. From the selected frames we compute an average RGB texture map $\overline{\mathbf{T}_{\mathbf{x}_{0}}}$ and combine it with the observed texture map $\mathbf{T}_{\mathbf{x}_{0}}$ from Eq. 2 to pass to the appearance encoder $\mathcal{E}_{\mathrm{a}}$. This average texture $\overline{\mathbf{T}_{\mathbf{x}_{0}}}$ serves as an appearance proxy in a single/two-view input setting.
Effects of this average texture are clearly visible in Fig. 7: pose-dependent clothing folds and wrinkles on the back of actors are only present in 2- and 4-view settings, whereas in the monocular setting, appearance on the back lacks pose-dependent effects, as they were not observed in the dynamic input stream.
Appendix E In-the-Wild Examples.
Neuman dataset [13] contains short monocular videos of a few subjects. We use Neural Localizer Fields (NLF) [47] to estimate SMPL-X parameters for actors in videos. We also compute the average texture for each actor, as described in Sec. D. Fig. 8 presents qualitative results produced by GIGA in the monocular input setting. Actors from Neuman are novel to GIGA, and GIGA’s training dataset did not contain any monocular videos. Despite that, renderings produced by GIGA are plausible and faithfully reconstruct observed actors.
Appendix F Additional Ablations.
@@
Scaling Dataset Size. GIGA also benefits from training on large data collections, as reported in Tab. 4 and Fig. 9. While scaling a digital human rendering model to hundreds of identities presents a challenge, it leads to an improved generalization overall.
@@
Original skeleton tracking and SMPL-X parameters in MVHumanNet are obtained from a dense set of views. To mimic a real-world scenario, we re-track a small number of sequences from a set of 4 views and fit SMPL-X models to estimated 3D keypoints. Qualitative results of GIGA with sparsely tracked SMPL-X template are shown in Fig. 10. If sparse tracking is mostly correctly aligned with the actual human actor, then GIGA can produce results, similar to dense view tracking case. For real-world use cases, a suitable single/sparse view motion tracking solution should be applied, for example NLF [47], as we did for Neuman dataset in Sec. E.